Circular No. 5750

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 H. U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik

(MPE); W. Lewin and E. Magnier, Massachusetts Institute of Technology; P. Predehl, G. Hasinger, W. Pietsch, B. Aschenbach, and J. Trümper, MPE; G. Fabbiano, Center for Astrophysics; J. van Paradijs, University of Amsterdam; L. Lubin, Princeton University; and R. Petre, Goddard Space Flight Center, communicate: “We report on the x-ray spectrum of SN 1993J obtained with ROSAT on Apr. 3.4 (IAUC 5747). In total, 212 counts were collected. The spectrum appears to be very flat, consistent with either a power law having a photon index of ∼ –0.7 or a thermal spectrum with kT > 4 keV. The NH column density is 3.4 × 10^{20} cm^{−2}, slightly below the galactic H I column density. This corresponds to a luminosity of 1.6 × 10^{39} erg/s in the range 0.1-2.4 keV for a distance of 3.25 Mpc. The optical/xray luminosity ratio is therefore ∼ 1000, similar to SN 1980K but in strong contrast to SN 1987A. This first x-ray observation of a supernova (whose progenitor was a red supergiant) near maximum optical light indicates a relatively hard spectrum in the ROSAT x-ray band. We plan to observe the source several times during the next few weeks.”

U. Munari, Asiago Astrophysical Observatory, corrects the photometry reported by Sostero et al. on IAUC 5741, using the magnitudes provided by Hanzl for HD 86677 (IAUC 5742): Mar. 31.972 UT, V = 10.94, B–V = +0.37. Further electronic photometry (Johnson system unless denoted Cousins by subscript ‘c’): Apr. 1.020 UT, V = 10.90, B–V = +0.40 (Munari); 1.660, V = 11.22 (M. Okyudo and T. Ishida, Nishiharima Observatory, Hyogo, Japan); 3.16, B_{c} = 12.21, B_{c}–R_{c} = +0.66 (P. Prugniel, Observatoire de Haute Provence); 5.00, B_{c} = 12.46, B_{c}–R_{c} = +0.76 (Prugniel); 5.489, V = 11.88 (T. Kato, Kyoto University); 5.549, B = 12.25 (Kato); 5.513, I_{c} = 11.24 (Kato); 5.614, V = 11.89 (Kato); 5.660, V = 11.98 (Okyudo and Ishida); 5.93, V = 11.87 ± 0.03, U–B = –0.47 ± 0.10, B–V = +0.50 ± 0.05, V–R = +0.52 ± 0.04, V–I = +0.38 ± 0.10 (M. Mikołajewski and K. Belczyński, N. Copernicus University, Toruń; reference star 24 UMa); 5.669, B = 12.25 (Kato); 5.713, I_{c} = 11.08: (Kato).

GRB 930309

Corrigendum. On IAUC 5749, the apparent position of the radio source was given; the J2000.0 position is α = 21^{h}32^{m}52^{s}.96, δ = +54^{o}40′15′′.94.

1993 April 6 Daniel W. E. Green